A 2 kg block is held in equilibrium on a incline of angle 60° by horizontal force, F aplied in the direction. if the coefficient between block and incline is 0.3. Determine
a) the minimum value of F
b) the normal force of the incline on the block
1
Expert's answer
2020-10-23T11:53:09-0400


a) According to the free body diagram and Newton's second law, find the magnitude of the force:


"Ox:f+F \text{cos}\theta−mg \text{sin}\theta=0\\
Oy:N−mg \text{cos}\theta−F\text{sin}\theta=0."

The force of friction is


"f=\mu N."

After substituting the force of friction, express the horizontal force. The force is

"F=\frac{mg(\text{sin}\theta-\mu\text{cos}\theta)}{\mu\text{sin}\theta+\text{cos}\theta}=18.5\text{ N}."

b) The normal force is


"N=mg \text{cos}\theta+F\text{sin}\theta=25.8\text{ N}."